A charged paint is spread in a very thin uniform layer over the surface of a plastic sphere of diameter 10.0 cm , giving it a charge of -10.0 μC .

a. Find the electric field just outside the paint layer.

b. Find the electric field 9.00 cm outside the surface of the paint layer.
Answer #1

Solution) Charge q = - 10 uC = - 10×10^(-6) C

Diameter d = 10 cm = 10×10^(-2) = 0.1 m

Radius r = d/2 = (0.1/2) = 0.05 m

(a) Electric field just outside the paint layer , |E| = ?

| E | = Kq/(r^2)

K = 9×10^(9) Nm^2/C^2

| E | = (9×10^(9))(10×10^(-6))/(0.05^2)

| E | = 36000×10^(3)

| E | = 36×10^(6) N/C

(b) r' = 9 cm = 9×10^(-2) = 0.09 m

d = r + r' = 0.05 + 0.09 = 0.14 m

| E | = Kq/(d^2)

| E | = (9×10^(9))(10×10^(-6))/(0.14^2)

| E | = 4591.83×10^(3)

| E | = 4.59×10^(6)

| E | = 4.6 × 10^(6) N/C